# GlucOS: Security, correctness, and simplicity for automated insulin delivery

### v2

@@ 8.2 Does GlucOS ensure pump operation while overcoming malicious pump drivers?
Age Group TIR without
biological invariant
Adults 49.93%
Adolescents 55.83%
Children 67.41%
Figure 13: Average proportion of time spent in range across 21 virtual humans when fluctuating their insulin sensitivity. The low values for TIR stress the importance of an end-to-end security mechanism, i.e., the biological invariant.
8.3 Is the biological invariant important?
We run experiments on the simulator to assess the risks of not considering our end-to-end security mechanism, i.e., the biological invariant. Figure 13 shows the average TIR across age groups. Across all virtual humans, we record an average TIR of 57.56%, even with our algorithmic security mechanism, when fluctuating insulin sensitivity. This corresponds to an average drop of 29.28% as a result of not considering the biological invariant. These results emphasize the importance of incorporating an end-to-end invariant and undermines results obtained from pure simulation. We present case-studies involving the biological invariant with a real user using GlucOS in Section 9.3.
9 Using GlucOS in the real world
@@
Our user study for the six students was approved by our university’s IRB, and Bob’s use GlucOS was determined by our university’s IRB to be “self evaluation” and exempt from IRB.
In our implementation, the BioKernel is an iOS app that consists of 5.1k lines of code in our trusted computing base. It communicates with a Dexcom G6 or G7 CGM and Omnipod DASH insulin pump via Bluetooth low energy, and runs the core closed-loop algorithm. For the BioKernel, we use the Swift programming language, the Actor abstraction for data-race-free concurrency, and JSON files stored on disk to capture persistent state. Although the BioKernel is a fully featured and stand alone automated insulin delivery system, GlucOS uses an event logging system that enables separate apps to implement complementary features while remaining isolated from the BioKernel.
9.1 Ethical considerations
In this section, we outline the steps we take to ensure that we uphold high ethical standards and ensure that the people using our system, are safe.
One of the co-authors of this paper is a board-certified Endocrinologist, who specializes in T1D. They are both a clinician with an active practice seeing patients living with T1D and a researcher who focuses on research on T1D. They designed the safety protocol, which defines the testing we conduct to ensure that the individuals using our system are safe, and the specific criteria we use to stop the study, if needed. Appendix A includes a more detailed discussion of our ethical considerations and safety protocol.
9.2 Is an ML-based closed-loop system safe and effective in practice?
@@
Glucose Range Reactive Reactive ML
range (mg/dl) label 67 days last wk. 1 wk
<54 L2 Hypo $0$%
54-69 L1 Hypo 0.59% 0.48% 0%
70-180 In range 96.35% 91.90% 97.22%
181-250 L1 Hyper 2.79% 7.62% 2.78%
>250 L2 Hyper 0.18% 0% 0%
Figure 14: The percent of time spent in different glucose ranges for the reactive safe model and the predictive ML model. Bold values represent the best results for each range.
Figure 14 shows the amount of time Bob spent in different glucose level ranges over the duration of our study. Data from the reactive safe model covers 67 days, followed by one week of data from our predictive ML model. We also show data for the last week of using the reactive safe model to show trends over time. Overall, Bob had the best results when using ML.
@@
9.3 Do biological invariant violations trigger in practice?
@@
9.4 Are predictive alerts useful for security?
@@
For the two month study, our participants decreased their average time in hypoglycemia from 2.7% down to 1.6%, and all individuals experienced a decrease in the amount of time spent in hypoglycemia. For context, participants in a state-of-the-art automated insulin delivery clinical trial experienced a decrease of time in hypoglycemia of 0.9% [15], showing that our 1.1% decrease is substantial.
@@
10 Related work
@@
Previous research has looked at the security of implanted medical devices in general [32, 16, 60], in addition to looking at insulin pumps in particular [39, 56], with more recent work looking at providing improved security [45, 9]. Also, recent work has looked at applying formal methods to insulin pumps for high assurance [55]. These works focus on devices and their communication channels. In contrast, with GlucOS we assume that these devices are correct and secure and focus our efforts on the software we use to run the automated insulin delivery system.
Researchers have started to look at combining untrustworthy ML with more traditional forms of automated decision making, like Markov Decision Processes, from a theoretical perspective [41, 50, 57, 20, 40]. These theoretical works focus on providing algorithms to prove desirable properties about a combined ML / Markov system, like robustness and stability. In contrast, we take advantage of domain knowledge and formulate a simple and practical system for one specific problem: automated insulin delivery.
11 Conclusion
@@
Our work improves the security of automated insulin delivery systems and also provides a framework for thinking about security in other critical healthcare technologies. Our exploration is an extreme example of personalized, distributed, and automated healthcare using trustworthy computer systems.
Appendix A Ethical and safety considerations
One individual, who we will call Bob, used GlucOS for 9 months. Six university students also used the predictive alerts from GlucOS for 2 months. We put in place a protocol where we would stop our study and ask people to stop using our system and revert back to their previous management regime. In our protocol, we have formal in-person meetings with Bob every week to check his data and make sure that he is not incurring risk as a result of using GlucOS, and weekly surveys for the six students, plus monthly meetings. We would have stopped our study if any of the following conditions occurred:
  • Excessive hypoglycemia. We define excessive hypoglycemia as having spent 8% or more of time during a week with CGM readings below 70 mg/dl or more than 1% below 54 mg/dl. These amounts of time spent in hypoglycemia would have represented a significant increase in time spent in hypoglycemia for participants.
  • Insufficient CGM data. If they had less than 70% of the time during a week with CGM readings, then the lack of CGM data would represent a decrease for the participants.
  • Insufficient closed-loop runs. We designed GlucOS to run closed-loop algorithms every five minutes, or 288 times per day. If Bob had a day where our closed-loop algorithm ran successfully less than 200 times, it would represent a fundamental flaw in the system.
  • Serious complications due to diabetes. Serious complications include any hospitalization for diabetes related issues, severe hypoglycemia where a participant was unable to recover from hypoglycemia themselves and had to get help from someone else, diabetic ketoacidosis, or seizure.
  • New medical diagnosis requiring significant attention. If anyone had received a new medical diagnosis during the study, the study would have had the potential to be a distraction for participants.
  • New mental health diagnosis. If anyone were struggling with mental health during the study, the study would have the potential to be a distraction.
In addition to our safety protocol, we followed best practices for ethical research to ensure that we conducted our study ethically.
  • IRB. Our study went through our university’s IRB process, where the study including six university students who used GlucOS’s alerting software was approved, and our case study on Bob was determined to be “self evaluation” and exempt from IRB.
  • Informed consent. All participants went through a written informed consent process.
  • Data management. GlucOS does not store any personally identifiable information and uses access control on our Google Cloud Platform infrastructure to ensure that only the authors of the paper can access this anonymized data.
  • Risk/benefit analysis. All participants are already living with T1D and use CGMs and inject insulin, so they already take on the risks inherent in managing T1D. The benefits include improved control over their T1D. There are risks in using experimental software, but all participants continued to rely on their CGM alerts as a last line of defense.
  • Participant selection. We selected participants who were university students living with T1D.
  • Adverse event reporting. Had any participants experienced adverse events during the study, we would have reported it to our IRB for guidance and addressed it according to our safety protocol.
  • Communication. All surveys were done through our university’s medical school infrastructure for private communication with individuals.
Since Bob used the full GlucOS software, including for insulin delivery, we included several other safety mechanisms:
  • Bob increased the frequency of his visits to his Endocrinologist while using the GlucOS software. He went from yearly appointments to quarterly appointments.
  • During the study, twice Bob got lab work based on blood tests to provide an independent measure of his overall metabolic health. This lab work helped to ensure that he was in good health and that there were no hidden negative impacts from his use of GlucOS.
Overall, GlucOS had a large positive impact on the individuals who used it. Bob’s lab-based A1C test (a blood test that measures average glucose levels over time), which includes his time using GlucOS, was 5.8%. That is nearly non-diabetic levels of control – healthy individuals will have an A1C of 5.6% or lower. All of the students decreased the amount of time spent in hypoglycemia. These results are profound for the seven people using GlucOS, and it had a positive impact on their health and quality of life.
Finally, Bob was a willing and enthusiastic user of the GlucOS system. To be clear, Bob is not a graduate student who was forced to use GlucOS so that his advisor could publish a paper. Rather, Bob struggled with the cognitive load of T1D management and wanted an automated insulin delivery system to offload part of the burden. However, he was unwilling to use any of the other automated insulin delivery systems out there due to concerns over their security. Bob found that GlucOS, with its focus on security, correctness, and simplicity, met his needs.
Appendix B List of pump setup and UX communication messages
Message Message Message
name type description
setupPod Setup Setup a new pump
assignAddress Setup Pair pump with phone
deactivatePod Setup Teardown pump
acknowledgeAlert UX Acknowledge alert from pump
configureAlerts UX Configure alerts from pump
beepConfig UX Configure pump beeps
Figure 15: Types of messages exchanged between a pump driver and pump pertaining to setup/UX.
We list insulin message types pertaining to pump setup and UX in Figure 15.
Figure 16: GlucOS’s CGM state machine to handle missing readings from CGMs.
Appendix C Runtime failure management
Figure 17: GlucOS’s state machine to transition users to the safest manual dosing state when losing trust in insulin delivery (i.e., with a buggy or malicious pump driver) or when the biological invariant is not satisfied.
In this section, we describe GlucOS’s state machine that combines its security components, i.e., algorithmic security (Section 3), driver security (Section 4) and end-to-end security (Section 5) together to protect users in light of their corresponding failures.
GlucOS’s state machine consists of three operating states: (1) ML with safety operating in closed loop, (2) reactive safe model operating in closed loop and (3) manual dosing in closed loop. ML provides the tightest level of control in managing glucose but poses the largest immediate threat from excess insulin since it preemptively injects insulin by anticipating rise in glucose. The reactive safe model provides slightly lesser control, but poses milder immediate threat since it only injects insulin to account for measured excess glucose. Manual dosing offers the least amount of control but poses no threat since it only injects the minimum insulin needed for sustenance. GlucOS predominantly operates in the state running ML (with safety) in closed loop.
We transition to the safest manual dosing state when we do not trust insulin delivery (i.e., with a buggy or malicious pump driver). We also transition to the safest manual operation when the biological invariant is not satisfied to ensure that we do not risk the possibility of removing too much glucose from our insulin doses. If we end up in the manual dosing state, we enforce that only manual user intervention can transition back to other states. We visualize these state transitions in Figure 17.
Inability to retrieve readings from a CGM also count as runtime failures. However, these failures are easier to handle.
Since ML needs a stream of CGM readings as input, we transition from ML to the reactive safe model if we don’t receive an input from the CGM. As long as the number of missing CGM readings is low (say, within 30 minutes), we can interpolate to account for the missing readings to operate the reactive safe model. Once we have too many missing CGM readings that make us lose confidence in interpolation, we transition to manual dosing. From manual dosing, we transition back to the reactive safe model on receiving a CGM reading. Lastly, from the reactive safe model, we transition back to ML once we have enough CGM readings that are required by the ML model. Figure 16 summarizes these transitions.
Bug Type Description
Division by Zero Error Potential division by zero error when calculating insulin dosage if the user entered the same start and end time for a basal rate, resulting in a duration of zero.
Negative Value Handling The insulin dosing calculation for microbolusing and basal rate erroneously allowed negative values.
Out-of-Range Value Handling The microbolusing function failed to handle cases where one of its calculation components was out of the expected range or resulted in a negative value.
Violation of Intended Postcondition The function responsible for setting the insulin delivery duration violated its intended postcondition by allowing the end date to be earlier than the start date.
Violation of Intended Postcondition The function responsible for calculating the total insulin dose violated its intended postcondition by returning an incorrect upper bound value under certain conditions.
Violation of Intended Postcondition The function responsible for updating the user’s insulin dosage violated its intended postcondition by failing to validate the input parameters correctly.
Violation of Intended Postcondition The function responsible for calculation of ‘unitsDelivered‘ needed additional postconditions accurately reflect the amount of insulin delivered over the time gap.
Lack of Error Handling Additional error handling was added to insulin delivered over a time segment function to gracefully catch undesirable state.
Lack of Error Handling Additional error handling was added to microbolusing function to gracefully catch undesirable states.
Figure 18: Report of bugs fixed by formally verifying GlucOS.
Appendix D Did formal verification help with correctness?
During the formal verification process of our secure insulin delivery system, we uncovered various bugs and inconsistencies that could have compromised the safety and reliability of the system in niche scenarios. We identified a total of 9 bugs across different categories. 3 of these bugs were related to invalid ranges for variables such as negative value returns, domain-specific out of range values, potential division by zero errors etc. in the insulin dosing calculations which were resolved by adding appropriate checks and error handling mechanisms to prevent such operations and ensure graceful handling of these cases. Additionally, we found 4 instances where the implemented functions violated their intended postconditions along the insulin delivery path, which were addressed by thoroughly reviewing and modifying the implementations to correctly satisfy the postconditions, preserving the intended behavior and meeting our claims for critical funcrions. Moreover, our analysis uncovered 2 situations where certain functions lacked proper error handling mechanisms, potentially leading to undefined behavior or system failures. By addressing these bugs and incorporating formal verification throughout the development lifecycle, we significantly enhanced the robustness, reliability, and safety of our secure insulin delivery system.
Appendix E Formal verification: invariants and bugs
In this section, we list the invariants used to formally verify critical functions of GlucOS. Figure 18 lists the bugs we found and fixed during formal verification.
E.1 Invariants
The MicroBolusing function ensures that micro boluses are always administered at safe time intervals while making sure that they are actually needed (the glucose level of the person is in fact high).
  • Ensure no micro bolus in the last 4.2 minutes.
  • Ensure glucose is at least 20 mg/dl above the target glucose level.
  • Ensure predicted glucose is greater than the current glucose level minus 2.
  • Ensure micro bolus amount is within the allowed range.
  • Ensure bolus amount is rounded correctly if pumpManager is available.
We assume that the BioKernel is the only software issuing bolus commands.
The GuardRails function is responsible for ensuring a safe basal rate value, mindful of the user’s glucose levels. It enforces several invariants:
  • Ensure the result of rounding newBasalRateRaw to the supported basal rate is within the allowable range.
  • Ensure the new basal rate does not exceed the maximum basal rate in the settings.
  • Ensure the new basal rate is not negative.
  • Ensure that if either the current glucose level or the predicted glucose level falls below or equals the shut-off glucose threshold, the new basal rate is set to 0.0.
The function operates under the assumption that all paths must go through the guard rails, with a specific evaluation comparing with the pump invariant. Additionally, it assumes the integrity of settings, trusting the underlying OS and file system and storage stack.
The insulinDeliveredForSegment function ensures that intersectionStart is not greater than intersectionEnd when computing the intersection between self.startDate and self.endDate with startDate and endDate.
  • Verifies that intersectionDuration is non-negative and does not exceed doseDuration, ensuring it accurately reflects the overlapping time duration.
  • Checks that the units of insulin (units) used for calculation are either deliveredUnits or programmedUnits, ensuring consistency in how the insulin amount is derived.
  • Verifies that the computation (units * intersectionDuration / doseDuration) accurately represents the amount of insulin delivered during the intersection period.
The function assumes that the input parameters (self.startDate, self.endDate, startDate, endDate, units, and doseDuration) are valid and within the expected ranges. It is also assumed that the underlying data structures and calculations are accurate and consistent throughout the system.
The CreateBasalDose function requires a time gap greater than 1 second to proceed and return a valid DoseEntry.
  • Ensures that the calculation of basalRatePerSecond from basalRate accurately represents the rate of insulin delivery per second.
  • Ensures that the calculation of unitsDelivered accurately reflects the amount of insulin delivered over the time gap.
  • The DoseEntry constructed should have consistent attributes and adhere to the specified type, units (unitsPerHour), and mutability (false).
The function assumes that the input parameters (basalRate, startDate, and endDate) are valid and within the expected ranges. It also assumes that the underlying time calculations and conversions between units are accurate and consistent.
The inferBasalDoses function ensures that basalDoses contains only DoseEntry instances where type is .tempBasal, .resume, or .suspend, sorted in ascending order by startDate.
  • InsulinType should be determined correctly based on the last .tempBasal or .bolus type dose in doses, defaulting to .humalog if none are found.
  • Each inferred basal dose added to inferredBasalDoses must be created using the createBasalDose function with valid parameters and added in the correct chronological order.
  • If the last dose in basalDoses is not of type .suspend, a valid basal dose must be inferred from its endDate to at.
The function assumes that the input parameters (doses, basalDoses, at, and pumpRecordsBasalProfileStartEvents) are valid and within the expected ranges. It also assumes the correctness of the underlying data structures and the createBasalDose function.
The insulinOnBoard function ensures that doses contains unique and valid DoseEntry instances filtered up to at.
  • The total iob should accurately represent the cumulative insulin on board from all valid dose entries in doses.
  • basalDoses should contain inferred basal doses that accurately reflect insulin on board contributions when pumpRecordsBasalProfileStartEvents is false.
  • The final return value of insulinOnBoard should be the sum of iob and basalIob, representing the total insulin on board at at.
The function assumes that the input parameters (doses, at, and pumpRecordsBasalProfileStartEvents) are valid and within the expected ranges. It also assumes the correctness of the underlying data structures and the inferBasalDoses function.
The ActiveBolus function ensures that data is read consistently from disk before proceeding with any operations.
  • Verifies that DeduplicatedDoses(events: eventLog, at: at) correctly filters out duplicates and retains only relevant dose entries up to at.
  • Ensures that doses.filter accurately selects bolus entries that were active at the specified at.
  • Verifies that doses.last correctly returns the last active bolus entry, ensuring it is non-null if a bolus exists within the specified criteria.
The function assumes that the input parameters (eventLog and at) are valid and within the expected ranges. It also assumes the correctness of the underlying data structures and the DeduplicatedDoses function, as well as the consistency and reliability of disk read operations.
The TempBasal (Safety Clamps) function ensures that the events array includes only events from safetyStates that fall within the time range [start - duration, at).
  • Verifies that the historicalMlInsulin value represents the total units of insulin delivered by the machine learning system over the specified time horizon.
  • Ensures that the machine learning and safety temp basal rates are correctly converted to units of insulin based on the given duration.
  • Ensures that the upper and lower bounds for insulin units are correctly adjusted based on historicalMlInsulin.
  • Ensures that the difference between machine learning and safety temp basal units is correctly clamped within the calculated bounds.
  • The clamped delta units should be converted back to a temp basal rate and added to the safety temp basal rate.
  • The returned SafetyTempBasal object should correctly encapsulate the adjusted temp basal rate and historical machine learning insulin.
The function assumes that the input parameters (safetyStates, start, duration, at, mlTempBasal, safetyTempBasal, and lastHistoricalMlInsulin) are valid and within the expected ranges. It also assumes the correctness of the underlying data structures and the conversion functions between units, as well as the proper configuration and appropriateness of the safety clamp parameters and thresholds for the user.

### v3

@@ 9.2. Does GlucOS ensure pump operation while overcoming malicious pump drivers?
10. Using GlucOS in the real world
@@
10.1. Ethical considerations
In this section, we outline the steps we take to ensure that we uphold high ethical standards and ensure that the people using our system, are safe. Our user study for the six students was approved by our university’s IRB, and Bob’s use GlucOS was determined by our university’s IRB to be “self evaluation” and exempt from IRB.
One of the co-authors of this paper is a board-certified Endocrinologist, who specializes in T1D. They designed the safety protocol, which defines the specific criteria we use to stop the study, if needed. The Appendix will include a more detailed discussion of our ethics and safety protocols.
10.2. Is an ML-based closed-loop system safe and effective in practice?
@@
Overall, Bob had the best results when using ML. His time spent in range increased from 96.3% to 97.2%, and notably he spent no time in hypoglycemia (below 70 mg/dl), down from 0.67% when not using ML.
@@
10.3. Do biological invariant violations trigger in practice?
@@
10.4. Are predictive alerts useful for security?
@@
For the two month study, our participants decreased their average time in hypoglycemia from 2.7% down to 1.6%, and all individuals experienced a decrease in the amount of time spent in hypoglycemia. For context, participants in a state-of-the-art automated insulin delivery clinical trial experienced a decrease of time in hypoglycemia of 0.9% (brown2019six, ), showing that our 1.1% decrease is substantial.
@@
10.5. How does the BioKernel perform?
We look at two aspects of system performance when evaluating our real-world use of GlucOS. First, we measure how long our closed-loop algorithm takes to run and second, we count how often the closed-loop algorithm runs successfully end-to-end. We use Bob’s iPhone 14 to collect all of the data we show here and these numbers are averages across the entire eleven month period.
On average, the BioKernel spends 4.3s running the closed loop algorithm, where almost half of that time is dominated by CGM and pump communications. Of the components within the closed-loop algorithm, the ML predictions take 2.2s on average and the safety logic takes 19ms (0.019s). Since the system runs in the background, we believe that iOS runs the device with reduced computational capacity to save power. But overall, 4.3s of wall-clock time for every five minutes our closed-loop algorithm runs with only a tiny fraction of that coming from our safety logic represents practical system performance.
For successful closed-loop end-to-end execution, over the evaluation period the closed-loop algorithm runs successfully 96.5% of the time, with 2.1% of the runs failing due to pump communication, which is our largest class of failure.
11. Related work
@@
The complexity of automated insulin delivery systems and the fact that they make real-time healthcare decisions, makes their security paramount. Current research on security for these systems focuses primarily on evaluating algorithms using simulations (zhou2021data, ; zhou2023knowsafe, ). We argue that this approach falls short of addressing the full spectrum of challenges inherent in building trustworthy systems that operate within real-world constraints. Real-world systems must cope with runtime platform limitations, changes in human physiology, and the necessity of human interaction with the system. Plus, algorithms and simulations model attacks abstractly through their glucose response, missing out on details like which components of the system are likely to be compromised or how the system architecture can prevent damage from these potentially vulnerable components. Our experience shows that when faced with these practical challenges, algorithms designed in simulations often prove insufficient or inapplicable. These real-world complexities require a fundamentally different approach to designing security for automated insulin delivery systems. To address these challenges, GlucOS moves beyond simulation-based designs and tackles the challenges of real-world systems.
Previous research has looked at the security of implanted medical devices in general (halperin2008security, ; burleson2012design, ; rushanan2014sok, ), in addition to looking at insulin pumps in particular (li2011hijacking, ; paul2011review, ), with more recent work looking at providing improved security (marin2016feasibility, ; ahmad2018securing, ). Also, recent work has looked at applying formal methods to insulin pumps for high assurance (panda2021secure, ). These works focus on devices and their communication channels. In contrast, with GlucOS we assume that these devices are correct and secure and focus our efforts on the software we use to run the automated insulin delivery system.
12. Conclusion
@@
Our exploration is an extreme example of personalized, distributed, and automated healthcare using trustworthy computer systems.